OpenStudy (anonymous):

Find the simplet polynomial with the following Zeros. -6 2i -2i

OpenStudy (anonymous):

f(x)=(x+6)(x-2i)(x+2i) f(x)=(x+6)(x^2-4i^2) f(x)=(x+6)(x^2+4) f(x)=x^3+6x^2+4x+24